You have a sample of gas that may be either N2O or toxic H2S. Testing the effusion of the sample and comparing it to the rate of effusion for N2, the unidentified gas effuses at 80% the rate of N2. What is the gas?
Answer #1

Answer: Nitrous Oxide is the gas that effuses at 80% the rate of N2.

Applying Graham's law of diffusion, we have

= Molecular weight of N2 / N2O

= 28/44= 0.636363

Root of ratio of molecular weight of N2   to N2O

=\sqrt{0.636363} = 0.79772

or 0.79772* 100 = 80%